OpenStudy (anonymous):

A ball is thrown across a playing field from a height of h = 5 ft above the ground at an angle of 45° to the horizontal at the speed of 20 ft/s. It can be deduced from physical principles that the path of the ball is modeled by the function y = − 32 (20)2 x2 + x + 5 where x is the distance in feet that the ball has traveled horizontally.

OpenStudy (anonymous):

dy/dx = -64x/400 + 1 = 0 x = 400/64 = 25/4

OpenStudy (anonymous):

now put 25/4 into the original equation

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

if it doesn't work show me what you did

OpenStudy (anonymous):

and i can walk you through

OpenStudy (anonymous):

oops i read hight of 6ft not 5

OpenStudy (anonymous):

oh ok i got it but now i need the horizontal distance ߺ

OpenStudy (anonymous):

a) differentiate y ' = 1 - (4/25) x max is at x such that y ' = 0 1 - (4/25) x = 0 → x = 25/4 after x = 6¼ ft the ball reaches its max height (substitute x = 25/4 in the formula which gives y) y = 8 1/8 ft (b) the ball hits the ground when its height is zero y = 0 (− 32/(20^2)) x^2 + x + 5 = 0 2x^2 - 25x - 125 = 0 x = (25 + √(25^2 - 4·2·(-125)))/4 ~ 16.3 ft

OpenStudy (anonymous):

b is the horizontal distance

OpenStudy (anonymous):

wait where does dx/dy come from in part 1?

OpenStudy (anonymous):

i mean dy/dx

OpenStudy (anonymous):

???

OpenStudy (anonymous):

i just the way i learned it

OpenStudy (anonymous):

its*

OpenStudy (anonymous):

its not important

OpenStudy (anonymous):

i mean how did you differentiate in part a?